Art calizon, jr. and batongbakal invested 80,000.00 and 40,000.00, respectively, in a hardware business. At the end oof the year , profits were 38,000.00. The partnership agreementprovided that each partner was to be credited with interest on his investment, and that the remaining profit (or loss) was to be shared ewually. If art received 2,400.00 more than the amount received by Orly, what was the rate of interest on the investment
Answer by mananth(12270)

let the rate of interest be x%
Orly invested 40000
so orly got an interest of 0.01x*40,000= 400x
Art received 2400 more than Orly
so he earned an interest of 400x +2400 on 80,000 at x%
..
400x+2400 = 0.01x*80,000
400x+2400 = 800x
800x-400x = 2400
400x=2400
x= 6%
...
m.ananth@hotmail.ca